OpenStudy (anonymous):

how is cosx=0 has general solution pi/2(2n+1)

OpenStudy (yttrium):

I think It is because x is initially pi/2. Therefore, using reference angle theorem, I think, we arrive at that solution.

OpenStudy (anonymous):

now my question id\s that if u consider that cosx=0 then cosx=cosx(pi/2) then by formula cos x=2npi+-y which in that case would be 2npi+-pi/2. then for any of sign either positive or negative how will u get this solution given above

OpenStudy (anonymous):

thanks in advance

OpenStudy (anonymous):

 cosx=0 when x=pi/2 and now by adding pi you will get other answers this means that general answer is \[\frac{ \pi }{ 2 }+n \pi=\frac{ \pi }{ 2 }+\frac{ 2n \pi }{ 2 }=\frac{ \pi }{ 2 }\left( 2n+1 \right)\]